# Quasi-steady evolution of fast neutrino-flavor conversions

### v1

@@ I Introduction
In dense neutrino gases, neutrino self-interaction [1, 2] can induce collective flavor conversions much faster than the neutrino oscillation in vacuum [3] and other relevant physical timescales in core-collapse supernovae (CCSNe) and binary neutron star mergers (BNSMs). The fast flavor conversion (FFC) represents such a case that has been attracting much attention in CCSN and BNSM modelers [4, 5, 6, 7, 8, 9, 10, 11]. Linear stability analysis of flavor conversions suggests that fast flavor instabilities (FFIs) commonly occur in CCSNe and BNSMs [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 14, 15, 12, 19, 20, 21], motivating detailed studies of the non-linear dynamics of FFCs [23, 24, 25, 26, 27, 28, 29, 30, 31, 32] including their interplay with various weak processes involving baryons and leptons [33, 34, 35, 36, 37, 38, 39, 40, 41, 36, 42].
@@
Very recently, on the other hand, the authors in [59] found an intriguing phenomenon, in which FFCs can persist even after spatially-averaged ELN-XLN angular crossings disappear if the background neutrinos are forced to be evolved (see Sec. III for more details). This results in different asymptotic states of neutrinos from those predicted by the canonical asymptotic schemes mentioned above (since flavor conversions are assumed to be no longer active after the disappearance of ELN-XLN angular crossings). It should be mentioned that the authors in [59] attempted to give an interpretation of the phenomenon with a quasi-linear theory, but it failed to reproduce outcomes, exhibiting that the mechanism is still far from being understood. Since this anomaly of FFC evolution could be a significant obstacle in developing accurate subgrid models, this problem is worthy of in-depth investigation. In this paper, we scrutinize the specific (but important) problem. As we shall show later, this analysis offers new insights into the non-linear properties of self-induced flavor conversions.
@@ V.2 Quasi-homogeneous analysis
As shown in Eq. 9, the increase rate of $\langle|\mathbf{P}_{R}|\rangle$ depends on not only $C_{R}$ but also $\langle P_{R,3}\rangle$. In SSSR model, $P_{R,3}$ (and $|\mathbf{P}_{R}|$) is zero at $t=0$ and then it increase with time with $\nu_{e}$ injection. In the very early phase (or linear phase of FFI), the norm grows linearly with time, $|\mathbf{P}_{R}|(t)=C_{R}t$, while the flavor coherence grows exponentially, $|\mathbf{P}^{T}_{R}|=\sqrt{P_{R,1}(t)^{2}+P_{R,2}(t)^{2}}\sim\mathrm{e}^{\mathrm{Im}[\omega(P_{R,3})]t}$, and henceforth the saturation is reached at smaller $|\mathbf{P}_{R}|$ (equivalently, smaller injection amounts) with lower $C_{R}$. Once the FFC enters the non-linear phase, $\mathbf{P}_{R}$ points in random directions in flavor space. This results in suppression of the increase of both $P_{R,3}$ and $|\mathbf{P}_{R}|$, which accounts for $|\mathbf{P}_{R}|\sim 0$ at $t<t_{m}$. Intuitively speaking, $\nu_{e}$ injection in the right-going neutrinos drives $\mathbf{P}_{L}$ dynamics so as to sustain $|\mathbf{P}_{R}|\sim 0$. As we shall show below, the flavor coherency for left-going neutrinos evolves with time, instead. On the other hand, $|\mathbf{P}_{R}|$ increases with time in the phase of $t>t_{m}$ until $\nu_{e}$ injection is ceased, which is because $P_{R,3}$ can no longer remain at zero; the detail will be discussed later.
In Fig. 4, the time evolution of the norm of the transverse component (or the first and second components) of $\mathbf{P}_{R/L}$ ($|\mathbf{P}^{T}_{L/R}|$) is also provided. The spatial average of $|\mathbf{P}^{T}_{L/R}|$ represents one of the key quantities that exhibits the overall degree of flavor coherency in the system. In fact, the degree of flavor coherency can be measured by $|\mathbf{P}^{T}_{L/R}|/|\mathbf{P}_{L/R}|$. As shown in the figure, $|\mathbf{P}^{T}_{L}|$ increases with time in the phase with $t<t_{m}$, and then it decreases afterward. This trend is consistent with our observation in $A_{L,3}$, which achieves the maximum flavor coherent state at $t=t_{m}$ ($A_{L,3}\sim 0$). It is worth noting that $|\mathbf{A}^{T}_{L/R}|$ is nearly zero during the entire simulation (which is also displayed in Fig. 4), exhibiting that the canonical homogeneous analysis does not have abilities to quantify the overall degree of flavor coherency.
@@
Figure 6: Schematic picture illustrating $\mathbf{P}_{L/R}$, $\beta_{L/R}$, and $\kappa$ under the quasi-homogeneous picture.
@@
The success of factorization indicates that the dynamics of $|\mathbf{P}_{L/R}|$ and the degree of flavor coherency (represented by $\beta$) are determined independently (or no correlations between the two variables). As we shall show below, $\langle\beta_{L/R}\rangle$ is a key variable to characterize the dynamics of the system.
As displayed in Fig. 5, $\langle\beta_{R}\rangle_{V}$ and $\langle\beta_{L}\rangle_{V}$ have remarkably different time evolution in the phase at $t<t_{m}$. The former remains nearly zero, whereas the latter linearly increases with time. These dynamical properties can be understood through the canonical picture of the non-linear phase of FFC. As already mentioned, right-going neutrinos experience strong flavor conversions to erase ELN-XLN angular crossings. As a result, $\mathbf{P}_{R}$ points in random directions, which makes $\langle\beta_{R}\rangle_{V}\sim 0$. For left-going neutrinos, on the other hand, the flavor coherency increases with time by FFC while preserving $|\mathbf{P}_{L}|$. In our quasi-homogeneous picture, the polarization vector rotates in flavor space (see also Fig. 6 as a schematic picture to see relations among key variables), and the time evolution of directional cosine with respect to the flavor axis is represented by $\langle\beta_{L}\rangle_{V}$.
One of the most striking findings in this study is that $\langle\beta_{L}\rangle$ and $\langle\beta_{R}\rangle$ are nearly equal to each other in the phase of $t>t_{m}$ for SSSR model (see Fig. 5), suggesting that the degrees of neutrino flavor polarization are correlated and balanced between the two neutrino beams. To see the correlation quantitatively, we compute the Kendall rank correlation coefficient concerning their spatial distributions. The time evolution of the coefficient is portrayed in Fig. 7. As shown in the figure, they are essentially no correlation until $t=t_{m}$ (which would be due to the nearly complete depolarization of $\mathbf{P}_{R}$), but they are in strong positive correlation at $t>t_{m}$. This indicates that the repolarization of $P_{R}$ is dictated by that of the left-going neutrinos.
@@
Figure 8: The same plots as in Fig. 2 and 5 but for SSSRs1 model.
@@ VI Semi-Analytic Model
Based on the results of our quasi-homogeneous analysis presented above, we develop a phenomenological model for the quasi-steady evolution of FFC found in SSSR model. We assume that right-going neutrinos are fully depolarized in the phase at $t<t_{m}$, i.e., $\langle|\mathbf{P}_{R}|\rangle_{V}=0$ and $\langle\beta_{R}\rangle_{V}=0$ (although one can analytically approximate its value from the linear phase). On the other hand, for left-going neutrinos, $\langle|\mathbf{P}_{L}|\rangle_{V}$ is constant in time. The time evolution of $\langle\beta_{L}\rangle_{V}$ can be obtained from the QKE for the zeroth angular moment of $\mathbf{A}_{0}(\equiv\mathbf{A}_{L}+\mathbf{A}_{R})$,
@@
which can be solved analytically. The solution can be expressed as
@@
We confirm that Eqs. 27 28 are in good agreement with numerical simulations (although numerical or statistical noises are present). One thing we learn from this analysis is that $\nu_{e}$ injection induces a correlation between $\langle\mathbf{P_{L}}\rangle_{V}$ and $\langle\mathbf{P_{R}}\rangle_{V}$, which dictates a quasi-steady evolution of the system. For a comprehensive understanding of the detail of mode coupling, we need more detailed studies, which will be discussed in forthcoming papers.
@@ VII Summary and discussion
    Our numerical simulations reproduced SR model in [59]. On the other hand, we found that the relation of $\langle A_{L/3}\rangle_{V}\sim\langle A_{R/3}\rangle_{V}$ appeared in the late phase is not an intrinsic property of quasi-steady evolution of FFC. In fact, the time evolution of $\langle A_{L,3}\rangle_{V}$ and $\langle A_{R,3}\rangle_{V}$ is remarkably different from each other for cases with slower neutrino injections (SSSR model).
@@
    Our phenomenological model also allows us to express the relation between $\langle\mathbf{P_{L}}\times\mathbf{P_{R}}\rangle_{V,3}$ and $C_{R}$ in SSSR model (see Eqs. 27 and 28). We note that $\langle\mathbf{P_{L}}\times\mathbf{P_{R}}\rangle_{V,3}$ corresponds to a convolution term of spatial wave number, which represents non-linear feedback from small-scale structures of flavor conversions. This provides a clue to study the correlation between $\mathbf{P_{L}}$ and $\mathbf{P_{R}}$, although the detailed investigation is postponed to future work.
@@ Acknowledgements.
H.N. is supported by Grant-in-Aid for Scientific Research (23K03468) and also by the NINS International Research Exchange Support Program. M.Z. is supported by Grant-in-Aid for JSPS Fellows (Grant No. 22KJ2906) and JSPS KAKENHI Grant Number JP24H02245. L.J. is supported by a Feynman Fellowship through LANL LDRD project number 20230788PRD1. R.A. is supported by JSPS Grants-in-Aid for Scientific Research (JP24K00632). S.Y. is supported by Grant-in-Aid for Scientific Research (21H01083). S.Y. is supported by the Institute for Advanced Theoretical and Experimental Physics, Waseda University, and the Waseda University Grant for Special Research Projects (project No. 2023C-141, 2024C-56, 2024Q-014). Numerical computation in this work was carried out at the Yukawa Institute Computer Facility. This work is also supported by the HPCI System Research Project (Project ID:240079).

### v2

@@ I Introduction
In dense neutrino gases, neutrino self-interaction [1, 2] can induce collective flavor conversions much faster than the neutrino oscillation in vacuum [3] and other relevant physical timescales in core-collapse supernovae (CCSNe) and binary neutron star mergers (BNSMs). The fast flavor conversion (FFC) represents such a case that has been attracting much attention in CCSN and BNSM modelers [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Linear stability analysis of flavor conversions suggests that fast flavor instabilities (FFIs) commonly occur in CCSNe and BNSMs [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 20, 21, 18, 25, 26, 27], motivating detailed studies of the non-linear dynamics of FFCs [29, 30, 31, 32, 33, 34, 35, 36, 37, 38] including their interplay with various weak processes involving baryons and leptons [39, 40, 41, 42, 43, 44, 45, 46, 47, 42, 48].
@@
Very recently, on the other hand, the authors in [61] found an intriguing phenomenon, in which FFCs can persist even after spatially-averaged ELN-XLN angular crossings disappear if the background neutrinos are forced to be evolved (see Sec. III for more details). This results in different asymptotic states of neutrinos from those predicted by the canonical asymptotic schemes mentioned above (since flavor conversions are assumed to be no longer active after the disappearance of ELN-XLN angular crossings). It should be mentioned that the authors in [61] attempted to give an interpretation of the observed FFC with a quasi-linear theory, but it did not apply to the novel case, exhibiting that the mechanism is still far from being understood. Since this anomaly of FFC evolution could be a significant obstacle in developing accurate subgrid models, this problem is worthy of in-depth investigation. In this paper, we scrutinize the specific (but important) problem. As we shall show later, this analysis offers new insights into the non-linear properties of self-induced flavor conversions.
@@ V.2 Quasi-homogeneous analysis
As shown in Eq. 9, the increase rate of $\langle|\mathbf{P}_{R}|\rangle$ depends on not only $C_{R}$ but also $\langle P_{R,3}\rangle$. In SSSR model, $P_{R,3}$ (and $|\mathbf{P}_{R}|$) is zero at $t=0$ and then it increase with time with $\nu_{e}$ injection. In the very early phase (or linear phase of FFI), the norm grows linearly with time, $|\mathbf{P}_{R}|(t)=C_{R}t$. Once the FFC enters the non-linear phase, $\mathbf{P}_{R}$ points in random directions in flavor space. This results in suppression of the increase of both $P_{R,3}$ and $|\mathbf{P}_{R}|$, which accounts for $|\mathbf{P}_{R}|\sim 0$ at $t<t_{m}$. Intuitively speaking, $\nu_{e}$ injection in the right-going neutrinos drives $\mathbf{P}_{L}$ dynamics so as to sustain $|\mathbf{P}_{R}|\sim 0$. As we shall show below, the flavor coherency for left-going neutrinos evolves with time, instead. On the other hand, $|\mathbf{P}_{R}|$ increases with time in the phase of $t>t_{m}$ until $\nu_{e}$ injection is ceased, which is because $P_{R,3}$ can no longer remain at zero; the detail will be discussed later.
In Fig. 4, the time evolution of the norm of the transverse component (or the first and second components) of $\mathbf{P}_{R/L}$ ($|\mathbf{P}^{T}_{L/R}|$) is also provided. The spatial average of $|\mathbf{P}^{T}_{L/R}|$ represents one of the key quantities that exhibits the overall degree of flavor coherency in the system. In fact, the degree of flavor coherency can be measured by $|\mathbf{P}^{T}_{L/R}|/|\mathbf{P}_{L/R}|$. As shown in the figure, $|\mathbf{P}^{T}_{L}|$ increases with time in the phase with $t<t_{m}$, and then it decreases afterward. This trend is consistent with our observation in $A_{L,3}$, which achieves the maximum flavor coherent state at $t=t_{m}$ (${\color[rgb]{0,0,0}\langle\beta_{L}\rangle_{V}}\sim 0$). It is worth noting that $|\mathbf{A}^{T}_{L/R}|$ is nearly zero during the entire simulation (which is also displayed in Fig. 4), exhibiting that the analysis based on only the homogeneous mode ($k=0$) is unable to quantify the overall degree of flavor coherency. The action of inhomogeneous modes ($k\neq 0$) is important for the evolution of FFC.
@@
Figure 6: Schematic picture illustrating $\mathbf{P}_{L/R}$, $\beta_{L/R}$, and $\kappa$ under the quasi-homogeneous picture. The picture is relevant for $t>t_{m}$ where $\langle\kappa\rangle_{V}>0$ is established.
@@
The success of factorization indicates that the dynamics of $|\mathbf{P}_{L/R}|$ and the degree of flavor coherency (represented by $\beta$) are determined independently. As we shall show below, $\langle\beta_{L/R}\rangle$ is a key variable to characterize the dynamics of the system.
As displayed in Fig. 5, $\langle\beta_{R}\rangle_{V}$ and $\langle\beta_{L}\rangle_{V}$ have remarkably different time evolution in the phase at $t<t_{m}$. The former remains nearly zero, whereas the latter linearly increases with time. These dynamical properties can be understood through the canonical picture of the non-linear phase of FFC. As already mentioned, right-going neutrinos experience strong flavor conversions to erase ELN-XLN angular crossings. As a result, $\mathbf{P}_{R}$ points in random directions, which makes $\langle\beta_{R}\rangle_{V}\sim 0$. For left-going neutrinos, on the other hand, the flavor coherency increases with time by FFC while preserving $|\mathbf{P}_{L}|$. In our quasi-homogeneous picture, the polarization vector rotates in flavor space (see also Fig. 6 as a schematic picture to see relations among key variables, which is relevant for $t>t_{m}$ where $\langle\kappa\rangle_{V}>0$ is established [^1] Notably, $\langle\kappa\rangle_{V}$ begins to peel away from 0 long before $t_{m}$. Note that for $t<t_{m}$, although $|\mathbf{P}_{R}|$ is small, it is finite. Therefore, the value of $\langle\kappa\rangle_{V}$ in the early phase is meaningful. Although the dynamics is not yet fully understood, it is related to the local synchronization of the azimuthal angle between $\mathbf{P_{L}}$ and $\mathbf{P}_{R}$ in the nonlinear regime. In fact, there can exist an attractor near the state where the average azimuthal angle agrees between the two beams. The detailed analysis is postponed to our future work.), and the time evolution of directional cosine with respect to the flavor axis is represented by $\langle\beta_{L}\rangle_{V}$.
One of the most striking findings in this study is that $\langle\beta_{L}\rangle$ and $\langle\beta_{R}\rangle$ are nearly equal to each other in the phase of $t>t_{m}$ for SSSR model (see Fig. 5), suggesting that the degrees of neutrino flavor polarization tend to get correlated on average between the left-going beam and the right-going beam. To see the correlation quantitatively, we compute the Kendall rank correlation coefficient concerning their spatial distributions. The time evolution of the coefficient is portrayed in Fig. 7. As shown in the figure, they are essentially no correlation until $t=t_{m}$ (which would be due to the nearly complete depolarization of $\mathbf{P}_{R}$), but they are in strong positive correlation at $t>t_{m}$. This indicates that the repolarization of $P_{R}$ is dictated by that of the left-going neutrinos.
Note that the correlation coefficient is not 1 even if $\langle\beta_{L}\rangle_{V}$ is nearly equal to $\langle\beta_{R}\rangle_{V}$. This can be understood from the schematic picture depicted in Fig. 6. $P_{R}$ rotates around $P_{L}$, indicating that $\beta_{L}$ and $\beta_{R}$ are in general different locally. On the other hand, as shown in Fig. 7, the correlation coefficient increase with time because $\kappa$ approaches 1. Furthermore, an increasing correlation coefficient confirms that $\beta_{L}$ and $\beta_{R}$ tend to be concordant at each spatial position, in addition to the co-evolution of $\langle\beta_{L}\rangle_{V}$ and $\langle\beta_{R}\rangle_{V}$.
@@
Figure 8: The same plots as in Fig. 5 but for SSSRs1 model.
@@ VI Semi-Analytic Model
Based on the results of our quasi-homogeneous analysis presented above, we develop a phenomenological model for the quasi-steady evolution of FFC found in SSSR model. We assume that right-going neutrinos are fully depolarized in the phase at $t<t_{m}$, i.e., $\langle|\mathbf{P}_{R}|\rangle_{V}=0$ and $\langle\beta_{R}\rangle_{V}=0$ (although one can analytically approximate its value from the linear phase [^2] In the linear phase, the z-component and norm grow linearly as $P_{R,3}(t)=|\mathbf{P}_{R}(t)|=C_{R}t$, while the flavor coherence grows exponentially, $|\mathbf{P}^{T}_{R}|=\sqrt{P_{R,1}(t)^{2}+P_{R,2}(t)^{2}}\sim\mathrm{e}^{\mathrm{Im}[\omega(P_{R,3})]t}$. The saturation of the linear phase is marked by the point at which the magnitude of flavor coherence has grown to be a great fraction of the norm of the polarization vector. Henceforth the saturation is reached at smaller $|\mathbf{P}_{R}|$ (equivalently, smaller injection amounts) with lower $C_{R}$. In the non-linear phase for $t<t_{m}$, $\mathbf{P}_{R}$ is depolarized in the xy plane, which makes $|\mathbf{P}_{R}|$ cease evolving.). On the other hand, for left-going neutrinos, $\langle|\mathbf{P}_{L}|\rangle_{V}$ is constant in time. The time evolution of $\langle\beta_{L}\rangle_{V}$ can be obtained from the QKE for the zeroth angular moment of $\mathbf{A}_{0}(\equiv\mathbf{A}_{L}+\mathbf{A}_{R})$,
@@
which can be solved analytically. In the analysis for SSSR model, $C_{L}$ is set to zero. The solution can be expressed as
@@
We confirm that Eqs. 27 28 are in good agreement with numerical simulations (although numerical or statistical noises are present). One thing we learn from this analysis is that $\nu_{e}$ injection induces a correlation between ${\color[rgb]{0,0,0}\beta}_{L}$ and ${\color[rgb]{0,0,0}\beta}_{R}$, which dictates a quasi-steady evolution of the system. For a comprehensive understanding of the detail of mode coupling, we need more detailed studies, which will be discussed in forthcoming papers.
@@ VII Summary and discussion
    Our numerical simulations reproduced SR model in [61]. On the other hand, we found that the relation of $A_{L{\color[rgb]{0,0,0},}3}\sim A_{R{\color[rgb]{0,0,0},}3}$ appeared in the late phase is not an intrinsic property of quasi-steady evolution of FFC. In fact, the time evolution of $A_{L,3}$ and $A_{R,3}$ is remarkably different from each other for cases with slower neutrino injections (SSSR model).
@@
    Our phenomenological model also allows us to express the relation between $\langle\mathbf{P_{L}}\times\mathbf{P_{R}}\rangle_{V,3}$ and $C_{R}$ in SSSR model (see Eqs. 27 and 28). We note that $\langle\mathbf{P_{L}}\times\mathbf{P_{R}}\rangle_{V,3}$ corresponds to a convolution term of spatial wave number, which represents non-linear feedback from small-scale structures of flavor conversions. This provides a clue to study the correlation between the two beams $\mathbf{P_{L}}$ and $\mathbf{P_{R}}$, although the detailed investigation is postponed to future work.
@@ Acknowledgements.
H.N. is supported by Grant-in-Aid for Scientific Research (23K03468) and also by the NINS International Research Exchange Support Program. M.Z. is supported by Grant-in-Aid for JSPS Fellows (Grant No. 22KJ2906) and JSPS KAKENHI Grant Number JP24H02245. L.J. is supported by a Feynman Fellowship through LANL LDRD project number 20230788PRD1. R.A. is supported by JSPS Grants-in-Aid for Scientific Research (JP24K00632). S.Y. is supported by Grant-in-Aid for Scientific Research (21H01083). S.Y. is supported by the Institute for Advanced Theoretical and Experimental Physics, Waseda University, and the Waseda University Grant for Special Research Projects (project No. 2023C-141, 2024C-56, 2024Q-014). Numerical computation in this work was carried out at the Yukawa Institute Computer Facility. This work is also supported by the HPCI System Research Project (Project ID:240079). We thank Damiano F.G. Fiorillo for valuable comments.